40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 4 16:43:50 2014, manasa, Update, General, IFO alignment update 
    Reply  Tue Feb 4 17:40:19 2014, manasa, Update, General, Ready for pump down?? 
       Reply  Tue Feb 4 23:01:24 2014, Jenne, Update, General, Ready for pump down?? 
          Reply  Wed Feb 5 15:26:57 2014, manasa, Update, General, Ready for pump down?? 
       Reply  Mon Feb 10 16:02:12 2014, Jenne, Update, General, PRM oplev clipped in vacuum Screenshot-Untitled_Window-1.png
          Reply  Mon Feb 10 16:20:28 2014, Jenne, Update, General, PRM QPD recentered 
             Reply  Mon Feb 10 18:59:25 2014, Jenne, Update, ASC, PRM ASC better, but not great yet 
                Reply  Mon Feb 10 19:56:10 2014, rana, Update, ASC, PRM ASC better, but not great yet 
Message ID: 9616     Entry time: Mon Feb 10 16:02:12 2014     In reply to: 9597     Reply to this: 9617
Author: Jenne 
Type: Update 
Category: General 
Subject: PRM oplev clipped in vacuum 

The PRC (locked on carrier so far today), is pretty wobbly.  It'll stay locked on carrier, but it's wobbling.  The ASC was over-ridden during the vent.  While I was looking around for that, I noticed that the PRM oplev sum is very low. 

I went into the lab, and turned off / blocked all oplev beams except the PRM beam.  I can't tell what it's clipping on, but there is definitely some red glow in the BS chamber (not as much as the stuff that's coming from the ITMY or SRM oplev hitting a tip tilt suspension - that giant spot went away when I turned off the ITMY/SRM oplev laser).  The beam going into the vacuum is nice and strong, but the beam coming out is very weak, and has a horizontal line of scatter through it, like it's clipped somewhere in pitch.  The PRM oplev sum is currently ~150 cts, when it should be closer to 2,000.

Screenshot-Untitled_Window-1.png

So far, this seems to be livable, but it's definitely disappointing. 

ELOG V3.1.3-